OpenStudy (anonymous):

if there is 1/6 of a cake and mitchell only wants 3/12 of it how much does he want?

OpenStudy (anonymous):

3/12 = 1/4 (1/4) (1/6) = 1/24

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np